OpenStudy (anonymous):

what are the real and complex solution of the polynomial equation x^3-64=0

OpenStudy (misty1212):

factor this one using the difference of two cubes \[a^3-b^3=(a-b)(a^2+ab+b^2)\]

OpenStudy (misty1212):

here you have \(x^3-64\) the same as \[x^3-4^3\] factors \[(x-4)(x^2+4x+16)\]

OpenStudy (misty1212):

\[x-4=0\\ x=4\] the other two you need to use the quadratic formula or complete the square for \[x^2+4x+16=0\]

OpenStudy (anonymous):

misty1212 would it be \[-2+2i \sqrt{3},-2-2i \sqrt{3}\]